Assume Fender produces only three guitars: the Stratocaster, Dreadnought and Telecaster. A limitation of 720 labor hours per week prevents Fender from meeting the sales demand for these products. Product information is as follows:

Stratocaster Dreadnought Telecaster
Unit selling price $768 $480 $1,008
Unit variable costs (480) (240) (864)
Unit contribution margin $288 $240 $144
Labor hours per unit 36 24 36

Required   
Determine the weekly contribution from each product when total labor hours are allocated to the product with the highest.
1. Unit selling price.
2. Unit contribution margin.
3. Contribution per labor hour.
(Hint: Each situation is independent of the others.)

4. Determine the opportunity cost the company will incur if management requires the weekly production of 20 Telecasters.
Answer #1

1. Highest selling price is of telecaster. So hours will be allocated to it.

Total hours = 720

Hours per unit = 36 hours

Total telecasters produced = 720/36 = 20 telecasters

Contribution = 20 telecasters * $144 = $2880

2. Highest contribution is of Stratocaster.

Hours per unit = 36

Number of units = 720/36 = 20 units

Weekly contribution = 20*$288 = $5760

3. Calculation of contribution per hour

Stratocaster = $288/36 hours = $8

Dreadnought = $240/24 = $10

Telecasters = $144/36 = $4

Since dreadnought has highest contribution per hour so all hours will be allocated to it.

Units produced = 720/24 = 30 units

Contribution= 30* $240 = $7200

4. If we have to produce 20 telecasters then

Contribution from these telecasters = 20*$144 = $2880

Remaining hours = 720 - (20*36 hours) = 0 hours

So if we produce 20 telecasters then we will not be able to produce dreadnought which has highest contribution per hour.

So contribution received from telecasters = $2880

Contribution lost from dreadnought = $7200

Opportunity cost = $7200 - $2800 = $4400

Feel free to ask any queries .

Also plz upvote it means a lot.. thank you
Answer #2

Assume Fender produces only three guitars: the Stratocaster, Dreadnought and Telecaster. A limitation of 720 labor hours per week prevents Fender from meeting the sales demand for these products. Product information is as follows:




StratocasterDreadnoughtTelecaster
Unit selling price

$1,056$660$1,386
Unit variable costs

(660)(330)(1,188)
Unit contribution margin

$396$330$198
Labor hours per unit

362436


Required   
Determine the weekly contribution from each product when total labor hours are allocated to the product with the highest.
1. Unit selling price.
2. Unit contribution margin.
3. Contribution per labor hour.
(Hint: Each situation is independent of the others.)



Determine the opportunity cost the company will incur if management requires the weekly production of 20 Telecasters. Hint: You want to maximize short-run profit. Think about which guitar is most profitable.

answered by: anonymous